I have the number 25. That was the total after taking 75% off. How do we find the original number? What is the mathematical equation?
Found 2 solutions by Earlsdon, stanbon:
Answer by Earlsdon(6294)

How about:
x-0.75x = 25 Where x is the original number!

Answer by stanbon(75887)

I have the number 25. That was the total after taking 75% off. How do we find the original number? What is the mathematical equation?
---------------
Let the original number be "x".
EQUATION:
x-0.25x = 25
0.75x = 25
x = 25/0.75
x = 33.3333....
Cheers,
Stan H.